Bill invested a total of $8000 in two accounts, one paying 10% interest and the other paying 15% interest. If the total interest earned was $1000, how much was invested at each rate?
Answer by mananth(16946)

Part I 10.00% per annum
Part II 15.00% per annum

x+y=8000 ------------------------1
10.00%x +15.00% y=$1,000.00
Multiply by 100
10x+15y =$100,000.00--------2

Multiply (1) by -10
we get

-10x-10y =-80000

Add this to (2)

5y= $20,000.00

divide by 5

y=$4,000.00 investment at 15%
Balance $4,000.00 investment at 10%